# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ Abstract
Fluorescence collection from individual emitters plays a key role in state detection and remote entanglement generation, fundamental functionalities in many quantum platforms. Planar photonics have been demonstrated for robust and scalable addressing of trapped-ion systems, motivating consideration of similar elements for the complementary challenge of photon collection. Here, using an argument from the reciprocity principle, we show that far-field photon collection efficiency can be simply expressed in terms of the fields associated with the collection optic at the emitter position alone. We calculate collection efficiencies into ideal paraxial and fully vectorial focused Gaussian modes parameterized in terms of focal waist, and further quantify the modest enhancements possible with more general beam profiles, establishing design requirements for efficient collection. Towards practical implementation, we design, fabricate, and characterize a diffractive collection element operating at $\lambda=397$ nm predicted to offer a total 0.25% collection efficiency into a single waveguide mode. A more efficient design with more demanding fabrication requirements would offer $1.14\%$, and we indicate avenues to improved devices approaching the limits predicted for ideal beams. We point out a particularly simple integrated waveguide configuration for polarization-based remote entanglement generation enabled by integrated collection.
@@
In the context of trapped-ion quantum information processing bruzewicz2019trapped , collection of light scattered from atomic ions provides a mechanism for qubit state readout nagourney1986shelved ; harty2014high , as well photonic entanglement of distant qubits simon2003robust ; moehring2007entanglement ; stephenson2020high . This functionality is typically implemented by means of a high numerical aperture (NA) objective collecting far-field radiation onto a detector. Pursuit of large-scale trap arrays, as well as compact and robust implementations of full experimental systems moses2023race , has motivated recent work towards integration of detector devices within ion trap chips for direct fluorescence capture todaro2021state ; setzer2021fluorescence ; reens2022high .
@@
Resonant cavities offer a route to order unity collection efficiency at rates beyond that set by free-space spontaneous emission CollectIonCavityMonroe2012 ; thompson2013coupling ; Kim2011CollectionFiber ; the associated technical challenges though, particularly for ions, motivate an understanding of limits to free-space collection more readily implemented in scalable platforms. Far-field collection efficiency in integrated settings is subject to the same NA limits as with bulk optics. Only preliminary estimates for collection efficiencies into planar photonic structures have been presented in the literature to date, with limited work towards optimal designs for integrated collection optics.
To precisely describe collection efficiencies achievable and establish efficient metrics for optimization, here we show that a simple argument from reciprocity allows expression of the polarization-dependent collection efficiency in terms of the electric field that would be radiated by the collection optic projected onto the radiating dipole, enabling straightforward quantification and optimization of photon collection with almost no approximations made. We analyze practically achievable values for idealized collection elements designed to couple to Gaussian beams, extending beyond the paraxial approximation as required for the high-NA focuses required for efficient collection. We present preliminary design and characterization of a planar single-mode photonic collection optic, demonstrating submicron waist focusing at $\lambda=397$ nm in agreement with designed performance, and with a predicted total 0.25% collection efficiency. An improved design with suppressed spurious sidelobe scatter is predicted to offer 1.14%. These considerations indicate the required field concentration is achievable in current designs at the blue/UV wavelengths of interest for most species knollmann2024integrated , and we discuss routes to more complete optimization to approach the limits predicted. We also point out a significant simplification possible for polarization-based remote entanglement generation using integrated collection into single-mode photonics and single-photon interference stephenson2019entanglement . Our work establishes a simple foundation on which to analyze/design collection elements, and lays a basis for more sophisticated optimal design.
@@
Figure 1: (a) Schematic depicting point dipole source $\mathbf{j}_{\mathrm{d}}$ corresponding to fluorescing ion and associated fields $\mathbf{E}_{\mathrm{d}}$ and $\mathbf{H}_{\mathrm{d}}$, along with fields radiated by the collection grating as if sourced by an effective current source $\mathbf{j}_{\mathrm{g}}$ launching amplitude in a single waveguide mode sourcing the grating. (b) Orthogonal polarization states radiated by an atom, defined with respect to an external, quantizing B-field, can be collected into identical modes of separate waveguides, e.g. quasi-TE.
The power coupling $\eta$ between these fields, i.e. the fraction of power radiated by the dipole collected into the waveguide mode, can be written as an overlap integral snyder1983optical
@@
where, with reference to Fig. 1(a), the $z=0$ plane is the plane just above the collection grating; however, the efficiency is independent of this choice. This expression can be explicitly evaluated for the dipole emission pattern and the field from the collection grating, as is done numerically below.
Lorentz reciprocity provides a simple and intuitive expression for $\eta$. Considering a (power normalized) source current associated with the emitting atom $\mathbf{j_{\mathrm{d}}}$ sourcing $\mathbf{E_{\mathrm{d}}}$ and $\mathbf{H_{\mathrm{d}}}$, and $\mathbf{j_{\mathrm{g}}}$ the effective source for the grating field $\mathbf{E_{\mathrm{g}}}$ and $\mathbf{H_{\mathrm{g}}}$, the fundamental reciprocity relation tells us that novotny2012principles
$\nabla\cdot\left(\mathbf{E_{\mathrm{d}}}\times\mathbf{H_{\mathrm{g}}^{*}}+\mathbf{E_{\mathrm{g}}^{*}}\times\mathbf{H_{\mathrm{d}}}\right)=-\mathbf{j_{\mathrm{d}}}\cdot\mathbf{E_{\mathrm{g}}^{*}}-\mathbf{j_{\mathrm{g}}^{*}}\cdot\mathbf{E_{\mathrm{d}}}$ (2)
We consider a radiating point dipole at $\mathbf{r}_{0}$ with polarization vector $\mathbf{p}=p_{0}\mathbf{\hat{p}}$ so that $\mathbf{j_{\mathrm{d}}}=-i\omega_{0}p_{0}\mathbf{p}\delta(\mathbf{r}-\mathbf{r_{0}})$, and $\mathbf{E}_{\mathrm{g}}$ the field of the beam emitted by the grating coupler with effective source lying below $z=0$. We can integrate over the infinite half-volume above $z=0$ and use the fact that in the far-field (on the hemisphere at $r\rightarrow\infty$), where the propagating fields are transverse, the quantity within the divergence above vanishes at all points pozar2011microwave to find
@@
$\eta=\frac{1}{16}\omega_{0}^{2}p_{0}^{2}\left\lvert\mathbf{\hat{p}}\cdot\mathbf{E}_{\mathrm{g}}^{*}(\mathbf{r}_{0})\right\rvert^{2},$ (4)
is expressed in terms of $\mathbf{E}_{\mathrm{g}}$ evaluated only at the ion location. This form also shows that the overlap integral (Eq. 1) is equivalent for any arbitrary plane or curved surface.
Expressing $\eta$ this way establishes a straightforward basis on which to design/optimize structures for collection into single modes, simply requiring appropriately polarized, tight focuses at the ion location to maximize the (normalized) field projected on the radiating dipole, $\left\lvert\mathbf{\hat{p}}\cdot\mathbf{E}_{\mathrm{g}}^{*}(\mathbf{r}_{0})\right\rvert$. For readout, we simply want to maximize total efficiency, whereas tailored polarization selectivity quantified by the vector dot product above is key for many schemes for generation of ion-photon and ion-ion entanglement as described above.
Figure 2: Paraxial and vectorial Gaussian beam coupling to dipole radiation at $\lambda=397$ nm. $\eta$ is calculated for $\pi$ and $\sigma_{\pm}$ radiation (Fig. 1) coupling to Gaussian beams with dominant linear polarization along $\mathbf{\hat{x}}$ and $\mathbf{\hat{y}}$, respectively. The same $\eta$ is predicted via Eq. 1 for the vectorial Gaussian’s overlap with dipole radiation in thegrating plane (red curve), and from the field projection of Eq. 4 (black curve). Overlap calculated via the paraxial approximation (blue) falls off from the exact result (red) for the small waists relevant to trapped ions. Dipole polarization $\sigma_{\pm}$ (dashed) couples half as well as $\pi$ (solid) by linearity (see main text) with the exception of the apertured power curves (green). The optimal beams (green), and the pi polarized Gaussian (solid red) approach 50% power coupling (top left) for infinitely tight focuses. A wavelength-waist Gaussian beam (red) performs 89% as well as the ideal beam (green) for $\pi$ polarized dipoles, and 80% optimally for $\sigma_{\pm}$. Corresponding to the apertured power (green crosses), the optimal beams profiles at the focus (sub plots) resemble a Gaussian spot both in shape and size, except, at the tightest focuses (A,B).
As a consistency check between both expressions for $\eta$ and to provide insight into required beam profiles, we numerically calculate $\eta$ according to both formulations. We calculate dipole emission coupling to ideal Gaussian modes represented both paraxially and fully vectorially, as well as the optimal beams for a given numerical aperture constraint. For simplicity, we set the grating beam $\mathbf{E_{\mathrm{g}}}$ to propagate along $z$, excluding a brief note about ideal beams from an angle.
@@
normalized to unit power in SI units. The Rayleigh range $z_{R}$, Gouy phase $\Psi(z)$, beam waist $w(z)=w_{0}\sqrt{1+(z/z_{R})^{2}}$, and radius of curvature $R(z)$ are as defined in SiegmanLasersBook . The blue lines in Fig. 2 show the overlap of these paraxial beams evaluated with $\pi$- and $\sigma_{\pm}$- polarized dipole radiation, assuming the beam focused at the ion location and for varying focal waist. Note that both $\sigma_{\pm}$ couple equally to the Gaussian due to its assumed linear polarization, and with half the $\eta$ of $\pi$ radiation since only the $\mathbf{\hat{y}}$-component is coupled. The increasing $\eta$ with decreasing waist is as expected due to the stronger $\mathbf{E}_{\mathrm{g}}$ at the ion location for tighter focuses. The apparent drop-off for wavelength-scale waists, however, arises due to a focal shift inherent to the breakdown of the paraxial approximation; the paraxial fields at the overlap plane correspond, if propagated exactly, to a focus before the ion.
@@
To obtain accurate $\eta$ values and verify the overlap integral for tight focuses of practical interest, we go beyond the paraxial approximation. To do so, we choose the particular solution to Maxwell’s equations most similar to the paraxial field profile at the focal plane of the ion by finding its angular-spectrum decomposition NanoOpticsBook ; wolf1959electromagnetic , then removing longitudinal polarizations from each plane-wave component. The analytic expression for this Fourier field (equation 12) is derived in appendix A. Phase evolution and subsequent inverse Fourier transform then gives the field profile at any other plane (appendix A).
@@
With this representation we calculate $\eta$ via the overlap integral (Eq. 1) and the field projection (Eq. 4). For the latter, we use the field of the vectorial Gaussian beam at the origin, rather than the paraxial, to be most accurate, although they are approximately consistent. The black and thick red lines in Figure 2 shows $\eta$ calculated both ways for fully vectorial focused beams, confirming $\eta$’s scaling with $w_{0}$ beyond the paraxial approximation’s validity. These results indicate wavelength-scale Gaussian waists provide $\eta$ values of multiple percent.
@@
We also explore the theoretical maximum performance of an optimal collection optic, which would mode-match the dipole emission over its aperture ($\mathbf{E}_{\mathrm{g}}=\mathbf{E}_{\mathrm{d}}$), thereby coupling the full incident power. For our Gaussian beams we define a relevant aperture as containing 99% of the power. For the field of Eq. 7 this corresponds to a radius $1.5w(z_{g})$, where $w(z_{g})$ is the beam waist in the grating plane at $z_{g}$. The green lines in Fig. 2 show the total incident radiation over this aperture. These curves demonstrate that a linearly-polarized Gaussian beam focused to $w_{0}=\lambda$ performs 89% as well as the ideal beam given the aperture (green) for $\pi$ polarized dipoles, and 80% optimally for $\sigma_{\pm}$. Notably, this is not true for large waists at the focus. There, the apertured power grows large simply due to a large $w(z_{g})$, however $\mathbf{E}_{\mathrm{g}}$ and $\mathbf{E}_{\mathrm{d}}$ are poorly mode-matched.
@@
We also compute the optimal field profiles at the ion plane corresponding to the ideal collection optic. Because $\eta$ is given by the field projection (eq. 4), this corresponds simply to the field maximizing the field at the ion location for a constrained aperture. To do this we simply set $\mathbf{E_{\mathrm{g}}}=\mathbf{E_{\mathrm{d}}}$ within the aperture (up to a power normalization factor), and 0 elsewhere, and propagate the field back to the ion location after applying time reversal (appendix A). We confirm this is the ideal field profile by verifying that the field projection (eq. 4) at the ion location matches the apertured power (green crosses in Fig. 2).
@@
The optimal intensity profiles are shown in the sub plots in Fig. 2). The beam appears qualitatively similar to a Gaussian beam for $w_{0}\gtrsim\lambda$, consistent with the Gaussian beam $\eta$ values (red) closely tracking the total apertured power (green). That continues up to impractically wide Gaussian beams ($\approx 2$ $\mu$m), because after that point the aperture on the grating plane begins to grow again. These visualizations may offer some intuition for strictly optimal field profile design for collection, though again we note our analysis shows for wavelength-scale beam waists the expected gain beyond focused Gaussians is modest.
@@
Figure 3: (a) SEM image of the fabricated grating coupler designed for 30 forward emission with submicron waist at 30 $\mu$m height above the chip surface. (b) and (c) Cross sections of simulated and measured beam profiles at $\lambda=397$ nm, respectively in the $x-z$ plane, showing multiple diffraction orders. (d) and (e) Cross section of the simulated and the measured beam profiles, respectively in the $x-y$ plane at the focus height at $z=30$ $\mu$m (within 1 $\mu$m height measurement uncertainty). (f) and (g) Beam profiles along transverse ($y$) and longitudinal ($x$) cuts, respectively, at the focus. Points are from the measurement, dashed line indicates FDTD simulation result. Red lines are Gaussian fit to the center lobe indicating $w_{y}=0.67$ and $w_{x}=0.78$ $\mu$m. (h) Simulated beam profile for single-order reverse-emitting grating.
@@
We frame the above discussion by spot size, in order to work independently from the exact position of the apperture. However, there are some details that do depend on the location $z_{g}$. While less general, the following still provides an approximate rule of thumb. Given a circular aperture $40$ $\mu$m below the ion of practical size ($\gg\lambda$), the optimal fill factor for the Gaussian beam is to have a radius that is 88.1% of the aperture size. This size trades off the power clipped at the aperture with the improved mode matching to maximize the field projection at the focus. We compute this by normalizing the beam power, before clipping the field to 0 outside the aperture.
@@
To compare performance achievable with planar optic collection elements with that predicted for an ideal target Gaussian beam, we design, fabricate and test a focusing collection grating intended for submicron spots at $\lambda=397$ nm and a focal height of 30 $\mu$m, using the approach to grating apodization, chirp, and line curvature described in beck2024grating . A single 80 nm-thick layer of $(\mathrm{Hf}\mathrm{O}_{2})_{x}(\mathrm{Al}_{2}\mathrm{O}_{3})_{1-x}$ composite HfO2Jaramillo23 with $n=1.967$ at $\lambda=397$ nm is used as the waveguide core. In the full stackup this core sits on 3 $\mu$m of thermal $\mathrm{Si}\mathrm{O}_{2}{}$ on Si substrates, with $1.5$ $\mu$m of top SiO2 cladding.
@@
Designs targeted a 30 average forward emission angle, corresponding to constructive interference considering the reflection from the Si substrate, and a focal waist along both dimensions of $0.5$ $\mu$m. A minimum feature size of 75 nm was used for the design. The average periodicity is ${\sim}350$ nm, with perturbation duty-cycles in the range of 0.2-0.4.
@@
Fig. 3(a) shows an SEM image of the fabricated device. Full 3D finite-difference-time-domain (FDTD) simulation of the structure (Fig. 3b and d) indicates emission into the targeted Gaussian beam at a focus at $z=30$ $\mu$m with a simulated waist of approximately $0.7$ $\mu$m. The fabricated grating was characterized in beam profiling measurements via emission imaging through a $\mathrm{NA}=0.95$ 50$\times$ objective. The good correspondence between experimentally measured emission and fit waists with the FDTD simulation (Fig. 3f,g) indicates that submicron waists as required for efficient collection are realizable with the design method used.
@@
Higher-order emission can be addressed most simply through emission at angles supporting only a single diffracted order mehta2017precise . An improved design emitting to the same focal waist and supporting only a single diffracted order at a $30^{\circ}$ reverse angle exhibits a FDTD-simulated total upwards efficiency of $56\%$ with significantly suppressed sidelobes (Fig. 3h). This results in a much enhanced predicted $\eta=1.14\%$, essentially lower than the values in Fig. 2 only due to radiation efficiency. The smaller periodicities required for diffraction at this angle result in smaller minimum feature sizes of 45 nm and somewhat more challenging fabrication.

### v3

@@ Abstract
Fluorescence collection from individual emitters plays a key role in state detection and remote entanglement generation, fundamental functionalities in many quantum platforms. Planar photonics have been demonstrated for robust and scalable addressing of trapped-ion systems, motivating consideration of similar elements for the complementary challenge of photon collection. Here, using an argument from the reciprocity principle, we show that far-field photon collection efficiency can be simply expressed in terms of the fields associated with the collection optic at the emitter position alone. We calculate collection efficiencies into ideal paraxial and fully vectorial focused Gaussian modes parameterized in terms of focal waist, and further quantify the modest enhancements possible with more general beam profiles, establishing design requirements for efficient collection. Towards practical implementation, we design, fabricate, and characterize two diffractive collection elements operating at $\lambda=397$ nm; a forward emitting design is predicted to offer 0.25% collection efficiency into a single waveguide mode, while a more efficient reverse-emitting design offers $1.14\%$ collection efficiency, albeit with more demanding fabrication requirements. Close agreement between simulated and measured emission for both designs indicates practicality of these collection efficiencies, and we indicate avenues to improved devices approaching the limits predicted for ideal beams. We point out a particularly simple integrated waveguide configuration for polarization-based remote entanglement generation enabled by integrated collection.
@@
In the context of trapped-ion quantum information processing bruzewicz2019trapped , collection of light scattered from atomic ions provides a mechanism for state measurement dehmelt1982monoion ; bergquist1986observation ; nagourney1986shelved ; sauter1986observation ; harty2014high , as well photonic entanglement of distant qubits simon2003robust ; moehring2007entanglement ; stephenson2020high ; krutyanskiy2023entanglement ; o2024fast ; saha2024high . This functionality is typically implemented by means of a high numerical aperture (NA) objective collecting far-field radiation onto a detector. Pursuit of large-scale trap arrays, as well as compact and robust implementations of full experimental systems moses2023race , has motivated recent work towards integration of detector devices within ion trap chips for direct fluorescence capture todaro2021state ; setzer2021fluorescence ; reens2022high .
@@
Resonant cavities offer a route to order unity collection efficiency at rates beyond that set by free-space spontaneous emission keller2004continuous ; wilk2007single ; CollectIonCavityMonroe2012 ; thompson2013coupling ; Kim2011CollectionFiber ; schupp2021interface ; krutyanskiy2023entanglement ; the associated technical challenges though, particularly for ions, motivate an understanding of limits to free-space collection more readily implemented in scalable platforms. Far-field collection efficiency in integrated settings is subject to the same NA limits as with bulk optics. Only preliminary estimates for collection efficiencies into planar photonic structures have been presented in the literature to date, with limited work towards optimal designs for integrated collection optics.
To precisely describe collection efficiencies achievable and establish efficient metrics for optimization, here we show that a simple argument from reciprocity allows expression of the polarization-dependent collection efficiency in terms of the electric field that would be radiated by the collection optic projected onto the radiating dipole, enabling straightforward quantification and optimization of photon collection with almost no approximations. We analyze practically achievable values for idealized collection elements designed to couple to Gaussian beams, extending beyond the paraxial approximation as required for the high-NA focuses required for efficient collection. We present preliminary design and characterization of planar single-mode photonic collection optics, demonstrating submicron waist focusing at $\lambda=397$ nm in agreement with designed performance, and with a predicted total 0.25% collection efficiency. An improved device with suppressed spurious sidelobe scatter is predicted to offer 1.14% collection efficiency. These considerations indicate the required field concentration is achievable in current designs at the blue/UV wavelengths of interest for most species knollmann2024integrated , and we discuss routes to more complete optimization to approach the limits predicted. We also point out a significant simplification possible for polarization-based remote entanglement generation using integrated collection into single-mode photonics and single-photon interference stephenson2019entanglement . Our work establishes a simple foundation on which to analyze/design collection elements, and lays a basis for more sophisticated optimal design.
@@
Figure 1: (a) Schematic depicting point dipole source $\mathbf{j}_{\mathrm{d}}$ corresponding to fluorescing ion and associated fields $\mathbf{E}_{\mathrm{d}}$ and $\mathbf{H}_{\mathrm{d}}$, along with fields radiated by the collection waveguide and grating as if sourced by an effective current source $\mathbf{j}_{\mathrm{g}}$ launching amplitude in the single waveguide mode sourcing the grating. Green/orange arrows (lines) represent power flows (intensity countours) associated with the fields labeled. The line at the origin defined as $z=z_{g}$ represents a choice of overlap integral (Eq. 1) evaluation plane near the grating. (b) Orthogonal polarization states radiated by an atom, defined with respect to an external, quantizing B-field, can be collected into identical modes of separate waveguides. Black features represent waveguides and gratings in the $x-y$ plane; double-sided arrows represent the dominant E-field polarization for either quasi-TE waveguide mode. For the orientation depicted, $\pi$ and $\sigma$-polarized light (with radiation patterns denoted by red and blue solid lines) is coupled into quasi-TE modes of separate waveguide channels.
The power coupling $\eta$ between these fields, i.e. the fraction of power radiated by the dipole collected into the waveguide mode, can be written as an overlap integral snyder1983optical between the power-normalized dipole and grating fields:
@@
For simplicity, we choose $z=z_{g}$, the plane just above the collection grating (Fig. 1a). However, the reasoning below shows $\eta$ is independent of this choice.
A simpler, equivalent expression for $\eta$ derives from Lorentz reciprocity. Considering a (power normalized) source current associated with the emitting atom $\mathbf{j_{\mathrm{d}}}$ sourcing $\mathbf{E_{\mathrm{d}}}$ and $\mathbf{H_{\mathrm{d}}}$, and $\mathbf{j_{\mathrm{g}}}$ the effective source for the grating field $\mathbf{E_{\mathrm{g}}}$ and $\mathbf{H_{\mathrm{g}}}$, the fundamental reciprocity relation tells us that novotny2012principles
$\nabla\cdot\left(\mathbf{E_{\mathrm{d}}}\times\mathbf{H_{\mathrm{g}}^{*}}+\mathbf{E_{\mathrm{g}}^{*}}\times\mathbf{H_{\mathrm{d}}}\right)=-\mathbf{j_{\mathrm{d}}}\cdot\mathbf{E_{\mathrm{g}}^{*}}-\mathbf{j_{\mathrm{g}}^{*}}\cdot\mathbf{E_{\mathrm{d}}}.$ (2)
We consider a radiating point dipole at $\mathbf{r}_{0}$ with polarization vector $\mathbf{p}=p_{0}\mathbf{\hat{p}}$ so that $\mathbf{j_{\mathrm{d}}}=-i\omega_{0}p_{0}\mathbf{p}\delta(\mathbf{r}-\mathbf{r_{0}})$, and $\mathbf{E}_{\mathrm{g}}$ the field of the beam emitted by the grating coupler with effective source lying below $z=z_{g}$. We can integrate over the infinite half-volume above $z=z_{g}$ and use the fact that in the far-field (on the hemisphere at $r\rightarrow\infty$, depicted in Fig. 1a), where the propagating fields are transverse, the quantity within the divergence above vanishes at all points pozar2011microwave to find
@@
$\eta=\frac{1}{16}\omega_{0}^{2}p_{0}^{2}\left\lvert\mathbf{\hat{p}}\cdot\mathbf{E}_{\mathrm{g}}^{*}(\mathbf{r}_{0})\right\rvert^{2}$ (4)
is expressed in terms of the power normalized $p_{0}$ and $\mathbf{E}_{\mathrm{g}}$ evaluated only at the ion location. This form also shows that the overlap integral (Eq. 1) is equivalent for any arbitrary plane or curved surface.
Expressing $\eta$ this way establishes a straightforward basis on which to design/optimize collection into single modes of any structure, simply requiring appropriately polarized, tight focuses at the ion location to maximize the (normalized) field projected on the radiating dipole, $\left\lvert\mathbf{\hat{p}}\cdot\mathbf{E}_{\mathrm{g}}^{*}(\mathbf{r}_{0})\right\rvert$. For readout, we simply maximize total efficiency, whereas tailored polarization selectivity quantified by the vector dot product above is key for many schemes for generation of ion-photon and ion-ion entanglement as described above. Additionally, positioning errors’ impact on $\eta$ is given by the spatial dependence of the grating field $\mathbf{E}_{\mathrm{g}}(\mathbf{r})$.
Figure 2: Paraxial and vectorial Gaussian beam coupling to dipole radiation at $\lambda=397$ nm. $\eta$ is calculated for $\pi$ (solid lines) and $\sigma_{\pm}$ (dashed lines) radiation (Fig. 1) coupling to Gaussian beams with dominant linear polarization along $\mathbf{\hat{x}}$ and $\mathbf{\hat{y}}$, respectively. Both methods of calculating efficiency $\eta$ (Eq. 1, 4) agree for exact vectorial fields (red curve) and are independent of ion height. Overlap calculated via the paraxial approximation (blue) falls off from the exact result (red) for the small waists relevant for high collection efficiency, shown here for 40 $\mu$m ion height. Dipole polarization $\sigma_{\pm}$ (dashed) couples half as well as $\pi$ (solid) by linearity (see main text) with the exception of the apertured power curves (black). The optimal beams (black), and the $\pi$ polarized Gaussian (solid red) approach 50% power coupling (top left) for infinitely tight focuses. A wavelength-waist Gaussian beam (red) performs 89% as well as the ideal beam (green) for $\pi$ polarized dipoles, and 80% optimally for $\sigma_{\pm}$. Corresponding to the apertured power (black dots), the optimal beams profiles in the focal plane are shown at right for the three labeled points A, B, and C in the left panel, and resemble a Gaussian spot both in shape and size, except at the tightest focuses (A,B).
As a consistency check between both expressions for $\eta$ and to provide insight into required beam profiles, we numerically calculate $\eta$ according to Eqs. 1 and 4. We calculate dipole emission coupling to ideal Gaussian modes represented both paraxially and fully vectorially, as well as the optimal beams for a given numerical aperture constraint. For simplicity, we set the grating beam emission angle to be vertical ($\theta=0$ in Fig. 1a), that is, propagating along $+z$ with $\mathbf{E_{\mathrm{g}}}$ primarily in the $x-y$ plane.
@@
normalized to unit power in SI units. The Rayleigh range $z_{R}$, Gouy phase $\Psi(z)$, beam waist $w(z)=w_{0}\sqrt{1+(z/z_{R})^{2}}$, and radius of curvature $R(z)$ are as defined in SiegmanLasersBook . The blue lines in Fig. 2 show the overlap of these paraxial beams evaluated with $\pi$- and $\sigma_{\pm}$- polarized dipole radiation, for a beam focused at the ion location and for varying focal waist. Note that both $\sigma_{\pm}$ couple equally to the Gaussian due to its assumed linear polarization, and with half the $\eta$ of $\pi$ radiation since only the $\mathbf{\hat{y}}$-component is coupled. The increasing $\eta$ with decreasing waist is as expected due to the stronger $\mathbf{E}_{\mathrm{g}}$ at the ion location for tighter focuses. The apparent drop-off for wavelength-scale waists, however, arises due to a focal shift inherent to the breakdown of the paraxial approximation; the paraxial fields at the overlap plane correspond, if propagated exactly, to a focus before the ion.
@@
To obtain the exact $\eta$ values and verify the overlap integral for tight focuses of practical interest, we go beyond the paraxial approximation. To do so, we choose the particular solution to Maxwell’s equations most similar to the paraxial field profile at the focal plane of the ion by finding its angular-spectrum decomposition NanoOpticsBook ; wolf1959electromagnetic , then removing longitudinal polarizations from each plane-wave component. The analytic expression for this Fourier field (equation 12) is derived in appendix A. Phase evolution and subsequent inverse Fourier transform then gives the field profile at any other plane (appendix A).
@@
With this fully vectorial focused beam, we calculate $\eta$ via the overlap integral (Eq. 1) and field projection (Eq. 4). Both calculations produce the same red lines in Fig. 2 , and confirm $\eta$’s scaling with $w_{0}$ beyond the paraxial approximation’s validity. These results indicate wavelength-scale Gaussian waists provide $\eta$ values of multiple percent. We note that calculating $\eta$ via field projection (Eq. 4) with the paraxial focal fields avoids the paraxial focal shift issue, and gives results within 40% of the exact values down to the tightest waists shown in Fig. 2.
@@
We also explore the theoretical maximum performance of an optimal collection optic, which would of course mode-match the dipole emission over its aperture ($\mathbf{E}_{\mathrm{g}}=\mathbf{E}_{\mathrm{d}}$), thereby coupling the full incident power. For our Gaussian beams we define a relevant aperture as containing 99% of the power. For the field of Eq. 7 this corresponds to a radius $1.5w(z_{g})$, where $w(z_{g})$ is the beam waist in the grating plane at $z_{g}$, which of course grows large for tight waists in the focal plane (consistent with numerical aperture). The black lines in Fig. 2 show the total incident radiation over this aperture. These curves demonstrate that a linearly-polarized Gaussian beam focused to $w_{0}=\lambda$ performs 89% as well as the ideal beam given the aperture (black) for $\pi$ polarized dipoles, and 80% optimally for $\sigma_{\pm}$. Note that the increase in apertured power visible in Fig. 2 for large $w_{0}$ is due to the fact that, in the limit of large focal waists, $w(z_{g})\sim w_{0}$ and hence the apertured power increases with $w_{0}$ despite poor mode matching between $\mathbf{E}_{\mathrm{g}}$ and $\mathbf{E}_{\mathrm{d}}$.
@@
We also compute the optimal field profiles at the ion plane corresponding to the ideal collection optic. Because $\eta$ is given by the field projection (Eq. 4), this corresponds simply to maximizing the field at the ion location for a constrained aperture. To do this we simply set $\mathbf{E_{\mathrm{g}}}=\mathbf{E_{\mathrm{d}}}$ within the aperture, and 0 elsewhere. We then renormalize the field to unit power, apply time reversal, and propagate the field back to the ion location (appendix A). We confirm that this is the ideal field profile by verifying that the field projection (Eq. 4) at the ion location matches the apertured power (labeled points in Fig. 2).
@@
The optimal intensity profiles are shown in the right panels in Fig. 2. The beam appears qualitatively similar to a Gaussian beam for $w_{0}\gtrsim\lambda$, consistent with the Gaussian beam $\eta$ values (red) closely tracking the total apertured power (black). That continues up to impractically wide Gaussian beams ($\approx 2$ $\mu$m), because after that point the aperture on the grating plane begins to grow again. These visualizations may offer some intuition for strictly optimal field profile design for collection, though again we note our analysis shows for wavelength-scale beam waists the expected gain beyond focused Gaussians is modest.
@@
Figure 3: (a) SEM image of the fabricated grating coupler designed for 30 forward emission with submicron waist at 30 $\mu$m height above the chip surface. Cross sections of simulated (b) and measured (c) beam profiles at $\lambda=397$ nm in the $x-z$ plane, showing multiple diffraction orders. (d) Cross section of the simulated (top) and the measured (bottom) beam profiles, respectively in the $x-y$ plane at the focus height at $z=30$ $\mu$m (within 1 $\mu$m height measurement uncertainty). (e) and (f) Beam profiles along transverse ($y$) and longitudinal ($x$) cuts, respectively, at the focus. Points are from the measurement, dashed line indicates FDTD simulation result. Red lines are Gaussian fit to the center lobe indicating $w_{y}=0.67$ and $w_{x}=0.78$ $\mu$m. Simulated (g) and measured (h) beam profile for single-order reverse-emitting grating. The spatial extent of the grating is indicated by the dashed gray line along the $x$ axis of (b), (c), (g), and (h). Measured data in (c) and (h) are obtained with 1 $\mu$m resolution along $z$.
@@
We frame the above discussion by spot size, in order to work independently from the position of the aperture. However, there are some details that do depend on the location $z_{g}$. While less general, the following still provides an approximate rule of thumb. Given a circular aperture $40$ $\mu$m below the ion and of practical size ($\gg\lambda$), the optimal fill factor for the Gaussian beam is to have a radius that is 88.1% of the aperture size. This size trades off the power clipped at the aperture with the improved mode matching to maximize the field projection at the focus. We compute this by normalizing the beam power, before clipping the field to 0 outside the aperture.
@@
To compare performance achievable with planar optic collection elements with that predicted for an ideal target Gaussian beam, we design, fabricate and test a focusing collection grating intended for sub-micron spots at $\lambda=397$ nm and a focal height of 30 $\mu$m, using the approach to grating apodization, chirp, and line curvature described in beck2024grating . A single 80 nm-thick layer of $(\mathrm{Hf}\mathrm{O}_{2})_{x}(\mathrm{Al}_{2}\mathrm{O}_{3})_{1-x}$ composite HfO2Jaramillo23 with $n=1.967$ at $\lambda=397$ nm is used as the waveguide core. In the full stackup this core sits on 3 $\mu$m of thermal $\mathrm{Si}\mathrm{O}_{2}{}$ on Si substrates, with $1.5$ $\mu$m of top SiO2 cladding.
@@
A first design targeted a 30 average forward emission angle, to produce constructive interference upon reflection from the Si substrate mehta2017precise for the thermal $\mathrm{Si}\mathrm{O}_{2}$ thickness employed here, and a focal waist along both dimensions of $0.5$ $\mu$m. A minimum feature size of 75 nm was used for the design. The average periodicity is ${\sim}350$ nm, with perturbation duty-cycles in the range of 0.2-0.4.
@@
Fig. 3(a) shows an SEM image of the fabricated device. Full 3D finite-difference-time-domain (FDTD) simulation of the structure (Fig. 3b and d, upper panel) indicates emission into the targeted Gaussian beam at a focus at $z=30$ $\mu$m with a simulated waist of approximately $0.7$ $\mu$m. The fabricated grating was characterized in beam profiling measurements via emission imaging through a $\mathrm{NA}=0.95$ 50$\times$ objective. The good correspondence between experimentally measured emission and fit waists with the FDTD simulation (Fig. 3e,f) indicates that submicron waists as required for efficient collection are realizable with the design method used.
@@
Higher-order emission can be addressed most simply through emission at angles supporting only a single diffracted order mehta2017precise . An improved design targeting the same focal waist in design and supporting only a single diffracted order at a $30^{\circ}$ reverse angle exhibits a FDTD-simulated total upwards efficiency of $56\%$ with significantly suppressed sidelobes (Fig. 3g). This results in a much enhanced predicted $\eta=1.14\%$ collection efficiency, essentially lower than the values in Fig. 2 only due to upwards radiation efficiency. The smaller periodicities required for diffraction at this angle result in smaller minimum feature sizes of 45 nm and somewhat more challenging fabrication, although roughly consistent with capabilities of advanced foundry photonics processes rakowski202045nm . The measured beam profile for a device fabricated to this design is shown in Fig. 3h. The fit beam waist of the measured grating emission is $w_{x}=0.82$ $\mu$m and $w_{y}=0.6$ $\mu$m, in agreement with the full FDTD-simulated waists of $w_{x}=0.82$ $\mu$m and $w_{y}=0.66$ $\mu$m.